OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 26 vehicles and 88 tires. How many motorcycles passed his house?

OpenStudy (anonymous):

c + m = 26 4c + 2m = 88 2c + 2m = 52 4c + 2m = 88 2c = 36 c = 18 m = 8 8 motorcycles